the curve y=x^2-x+1 is translated by (0,3)
find and simplify the equation of the translated curve
1
Expert's answer
2021-02-24T07:33:15-0500

1) If the graph of  "y=f(x)"  is translated "a" units horizontally and "b" units vertically, then the equation of the translated graph is  "y-b=f(x-a)."



"y-0=(x-2)^2+3(x-2)+4""y=x^2-4x+4+3x-6+4""y=x^2-x+2"

2) We start with the graph of "y=f(x)." The graph of "g(x)=f(-x)" is the graph of "y=f(x)" reflected in the "y" -axis . How points in graph of "f(x)" become points of new graph "g(x)" ?



"(a,b)\to(-a,b)"

Visual effect: flip over the "y" -axis

The graph of "h(x)=3g(x)=3f(-x)" is the graph of "y=g(x)" stretched vertically by "3" units.

How points in graph of "g(x)" become points of new graph "h(x)"?



"(-a,b)\to(-a,3b)"

Visual effect: stretch vertically by "3."


How points in graph of "y=f(x)" become points of new graph "y=3f(-x)"?



"(a,b)\to(-a,3b)"